OpenStudy (anonymous):

Solve for x. - 4(x + 3) = 20

OpenStudy (anonymous):

-8

OpenStudy (shamim):

x+3=-20/4

OpenStudy (anonymous):

it becomes -4x -12 =20 therefore -4x=32 therefore x=-8

OpenStudy (anonymous):

(x+3) = 20/-4 x+3 = -5 x-3 = -5 -3 x = -8